Evaluate the function t = 83
R(83) = 0.01 (83) + 46.4
= 0.01(83) + 46.4
= 47.23
The correct answer is 45.57 How did they get the answer I need step by step help
45.52 = 0.01t + 46.4
-09 = 0.01 Where did they get the -0.9 and how what did they do
= 88.0 How did they get 88.0 also

Answer by stanbon(75887)

Evaluate the function t = 83
R(83) = 0.01 (83) + 46.4
= 0.01(83) + 46.4
= 47.23
The correct answer is 45.57 How did they get the answer I need step by step help
45.52 = 0.01t + 46.4
-09 = 0.01 Where did they get the -0.9 and how what did they do
= 88.0 How did they get 88.0 also
=======================================
The work you have shown is correct if R(t) = 0.01t+46.3
----
Cheers,
Stan H.